OpenStudy (anonymous):

points (0,0) and (3,3) lie on line k. what is the slope of the line that is perpendicular to k? A.1 B.-1 C.0 D.Undefined

OpenStudy (anonymous):

the answer is A. \[slope between two points = (y_2 - y_1)/(x_2-x_1)\] \[(3 - 0)/(3-0)\] \[3/3\]

OpenStudy (anonymous):

oh, perpendicular. My apologies. The slope of the line is 1. however the slope of its perpendicular is -1.

OpenStudy (anonymous):

the production of slope of the perpendikular lines equal to -1(minus 1). slope of the line k=(3-0)/(3-0)=3/3=1 so that the slope of the line perpendicular to k= -1

OpenStudy (anonymous):

The line that passes through two points is.. \[\LARGE y-y_1=\underbrace{\boxed{\frac{y_2-y_1}{x_2-x_1}}}_{k}(x-x_1)\] prependicular is : \[\LARGE k\cdot k_1=-1\] so we have: \[\LARGE k=\frac33\] hence: \[\LARGE \frac33\cdot k_1=-1\] \[\LARGE 1\cdot k_1=-1\] \[\LARGE k_1=-1\]